# Groups Acting Acylindrically on Trees

### v5

@@ 4.2.1. $I$-bundles
We will require some results about a special type of 3-manifold that is geometrically “nice” in terms of its fundamental group. We will briefly define the concept of an $I$-bundle over a surface, and include a strong result pertaining to $3$-manifolds whose boundary components induce finite index subgroups of the fundamental group. For proof and further exposition see Chapter 10 of [Hempel76].
@@
[Hempel76, Theorem 10.5] Let $M$ be a compact 3-manifold which contains no 2-sided real projective plane and let $S$ be a compact, connected, incompressible surface embedded in $\partial M$ such that $S$ is not homeomorphic to the disc $B^{2}$, the sphere $S^{2}$ or the projective plane $P^{2}$. If the index $[\pi_{1}(M)\colon i_{*}\pi_{1}(S)]$ is finite then either:
@@
    $[\pi_{1}(M)\colon\pi_{1}(S)]=1$ and $M=S\times I$ with $S=S\times\{0\}$;
@@
    $[\pi_{1}(M)\colon\pi_{1}(S)]=2$ and $M$ is an $I$-bundle over a compact surface $\overline{S}$, with $S$ a two sheeted cover of $\overline{S}$. We call such an $I$-bundle twisted.
@@
All manifolds we consider in this paper will be orientable, and thus cannot contain 2-sided real projective planes due to the fact that, if $P^{2}$ is the real projective plane, $P^{2}\times[-1,1]$ is non-orientable. The condition that $M$ has no two sided $P^{2}$ has however been included in the statement of Theorem 4.14 for completeness.
@@ 4.3. A Geometric Condition for Acylindrical Arboreality
Let $H$ be a non-trivial virtual surface fibre subgroup of $G$. Then by definition $H$ is finitely generated, and there exists $t\in G$ such that $F=H\rtimes\langle t\rangle$ is a finite index subgroup of $G$ corresponding to a finite sheeted cover of $M$ in which $H$ is a surface fibre subgroup. Let $A$ be subgroup of $G$ that contains $H$. We separate into two cases.
The first case occurs when $F\cap A=H$. In this case, $H$ has finite index in $A$ and is finitely generated, implying that $A$ is also finitely generated. Assume for contradiction that $A$ is not a virtual surface fibre subgroup of $G$. It then follows by Theorem 4.12 that $A$ must be geometrically finite as it is finitely generated, and so by Theorems 4.11 and 4.2 we have that $A$ has finite relative height in $G$ with respect to the boundary subgroups of $G$, which are trivial as $M$ is closed. For all $0\neq j\in\mathbb{Z}$ we have that $t^{j}\notin A$, so the set of conjugates $\{t^{j}At^{-j}\}_{j\in\mathbb{\mathbb{Z}}}$ is essentially disjoint, and the intersection of all of these subgroup will contain $H$ as every power of $t$ fixes $H$ by conjugation. Thus, $H$ must be a peripheral or finite subgroup of $G$ by the fact that $A$ has finite relative height, and all of our peripheral subgroups are trivial in this case, so $H$ must be finite. The manifold $M$ is orientable with empty boundary and irreducible by Theorem 4.8, and $G$ admits a non-trivial splitting so it must be infinite, so we may apply Lemma 4.10 to see that $G$ is torsion free, and $H$ in this case must be trivial. However, by Corollary 4.9, $\pi_{1}(M)$ admits a non-trivial splitting over the free subgroup if and only if $M$ is a solid torus, which is contradicts that $M$ is closed. It follows that that $H$ cannot be trivial, leading to the desired contradiction. Thus, in the case where $A\cap F=H$ we must have that $A$ is a virtual surface fibre subgroup.
The second case occurs when $A\cap F$ contains $H$ as a proper subgroup. In this case there exists $0\neq l\in\mathbb{Z}$ and $h\in H$ such that $t^{l}h\in A\cap F$, but $h\in A\cap F$ by assumption that $A$ contains $H$, and so we must have that $t^{l}\in A\cap F$. It follows that $A\cap F$ contains the subgroup $\langle t^{l},H\rangle$, which has finite index in $H\rtimes\langle t\rangle$, which in turn has finite index in $G$. Thus in this case $A$ contains the finite index subgroup $\langle t^{l},H\rangle$ of $G$, and so $A$ must have finite index in $G$. Thus in both cases either $A$ is a virtual surface fibre subgroup of $G$ or $A$ has finite index in $G$, as required. ∎
@@
First assume that $\Gamma$ has $2$ vertices, so we can express $G$ as a non-trivial amalgam $A*_{H}B$ of two subgroups $A$ and $B$ along $H$. In this case, we claim that the subgroup $H$ must have finite index in both $A$ and $B$. Indeed, we first observe that as the amalgam decomposition $A*_{H}B$ is non-trivial by hypothesis on $H$, neither $A$ nor $B$ can have finite index in $G$. Therefore both $A$ and $B$ must be virtual surface fibre subgroups of $G$ by Lemma 4.15, so the normalisers $N_{G}(A)$ and $N_{G}(B)$ of $A$ and $B$ in $G$ must have finite index in $G$. By elementary properties of amalgamated free products and non-triviality of the amalgam $A*_{H}B$, the normaliser $N_{G}(A)$ of $A$ in $G$ must have that $N_{G}(A)\cap B=H$, so it follows from the fact that $N_{G}(A)$ has finite index in $G$ that $H=N_{G}(A)\cap B$ must also be finite index in $B$. Similarly $H$ will have finite index in $A$, and the claim is proved.
The subgroup $H$ is a virtual surface fibre subgroup so it must be isomorphic to the fundamental group of a closed surface, and $M$ is irreducible by Theorem 4.8, so we may apply Theorem 4.5 to realise $H$ geometrically as the fundamental group of some incompressible surface $\Sigma$ in $M$. The manifold $M$ cut along $\Sigma$ then has two components $M_{1}$, with $\pi_{1}(M_{1})=A$, and $M_{2}$, with $\pi_{1}(M_{2})=B$, both 3-manifolds whose fundamental groups are virtually $\pi_{1}(\Sigma)$. In particular, $\Sigma$ must be $2$-sided as it separates $M$, so by the remark after Theorem 4.14 $\Sigma$ is not a real projective plane, and more generally neither $M_{1}$ nor $M_{2}$ can contain a $2$-sided projective plane by the same remark. Furthermore, $\Sigma$ is not a disc as $M$ is closed, and $\Sigma$ is not a sphere by Theorem 4.8. Thus, using again the fact that $M$ is irreducible and so each of $M_{1}$ and $M_{2}$ must be irreducible, we may invoke theorem 4.14 on $M_{1}$ (and respectively on $M_{2}$) to see that one of the following occurs.
@@
    $[\pi_{1}(M_{1}):\pi_{1}(\Sigma)]=1$, which cannot occur as our original splitting was non-trivial.
@@
    $[\pi_{1}(M_{1}):\pi_{1}(\Sigma)]=2$ and $M_{1}$ is a twisted $I$-bundle with boundary $\Sigma$.
@@
For the other case assume that $\Gamma$ has exactly $1$ vertex, which we will label $v$. We can then express $G$ as an $HNN$ extension of $G_{v}$ along $H$, so call the stable letter of this extension $\gamma$. We claim that in this case the subgroup $G_{v}$ must be equal to $H$, and must be a genuine surface fibre subgroup of $G$, i.e. there exists some fibration of $M$ over the circle whose fibre is a surface with fundamental group $G_{v}=H$. Indeed, we must have that either $G_{v}$ is either finite index in $G$ or that $G_{v}$ is a virtual surface fibre subgroup of $G$ by Lemma 4.15. Since $G$ is an HNN-extension with base group $G_{v}$, the subgroup $G_{v}$ cannot have finite index, and so we must in fact have that $G_{v}$ is a virtual surface fibre subgroup of $G$. It follows that there exists $t_{v}\in G$ such that $F_{v}=G_{v}\rtimes\langle t_{v}\rangle$ is a finite index subgroup of $G$ corresponding to a cover of $M$ in which $G_{v}$ is a surface fibre subgroup.
Thus there exists some integer $l>0$ such that $\gamma^{l}\in F_{v}$, and so both $\gamma^{l}$ and $\gamma^{-l}$ normalise $G_{v}$. However, by definition of $\gamma$ as the stable letter in the HNN-extension of $G_{v}$ along $H$ and by elementary properties of HNN-extensions of groups, this can only occur if $H=G_{v}$. It follows that $G=H\rtimes\langle\gamma\rangle$, $H$ is finitely generated by definition of a surface fibre subgroup, and $H\not\cong\mathbb{Z}/2\mathbb{Z}$ owing to the fact that by Lemma 4.10 and the fact that $G$ admits a non-trivial splitting and is therefore infinite we have $H\leq G$ is torsion free, so we may invoke Theorem 4.4 (algebraic fibring) to see that that $H=G_{v}$ is a surface fibre subgroup of $G$. The claim is therefore proved, and so in this case $H$ is a surface fibre subgroup so is normal by definition as required.
@@
For the if direction assume that $M$ contains an embedded $2$-sided incompressible geometrically finite closed surface $\Sigma$ that is not isotopic to any boundary component of $M$. Then $\Sigma$ is not a disc as it is closed, $\Sigma$ is not a projective plane by the remark following Theorem 4.14, $\Sigma$ is not a sphere by Theorem 4.8 and $\pi_{1}(\Sigma)$ is finitely generated as $\Sigma$ is closed. Since $\Sigma$ is geometrically finite we have that $H=\pi_{1}(\Sigma)$ is relatively quasi-convex by Theorem 4.11, and so by Theorem 4.2 there exists a natural number $n_{H}$ such that the intersection of $n_{H}$ essentially disjoint conjugates of $H$ in $G$ cannot contain any hyperbolic elements. In fact, it follows directly from the definition of relatively quasi-convex and by considering the action of $G$ on $\partial\mathbb{H}^{3}$ that $H$ is hyperbolic relative to the subgroups that arise as the intersections of $\Sigma$ with $\partial M$, and that all peripheral elements of $G$ that are included in $H$ are conjugate in $H$ to some boundary component of $S$. Thus, since $\Sigma$ is closed and therefore does not intersect with $\partial M$, $H$ contains only hyperbolic elements and so the intersection of any collection of at least $n_{H}$ essentially disjoint conjugates of $H$ in $G$ must be the trivial subgroup.
@@
It follows that the splitting $(\Gamma,\mathfrak{G})$ of $G$ induced by cutting $M$ along $\Sigma$ induces an $(n_{H},1)$-acylindrical action on its Bass-Serre tree $T$. The group $G$ is not virtually cyclic as it contains the fundamental group of $\Sigma$, a closed surface that is not a sphere or a projective plane. Thus the action of $G$ on $T$ is not lineal by Theorem 2.3, and to show that it is non-elementary it only remains to check that it is non-trivial.
@@
Assume for contradiction that this splitting is trivial, so for some vertex $v\in V(\Gamma)$, $G_{v}=G$. The graph $\Gamma$ has one edge, which corresponds to $\Sigma$, and so $\Gamma$ either has one vertex or two vertices. The former of these cannot occur, else $G$ would be an HNN-extension of $G_{v}$, so $G_{v}\neq G$. We may thus assume that the splitting $(\Gamma,\mathfrak{G})$ is an amalgam decomposition $G\cong G_{v}*_{H}A$ for some subgroup $A$, with $H=A$. Since $\Sigma$ is $2$-sided it must locally separate $M$, and so since the cut along $\Sigma$ induces an amalgam we must have that $\Sigma$ separates $M$ into two components $M_{1}$ and $M_{2}$ with $A=\pi_{1}(M_{1})$ and $G_{v}=\pi_{1}(M_{2})$. By assumption we have that $\pi_{1}(\Sigma)=\pi_{1}(M_{1})$, $\Sigma$ is a connected component of $\partial M_{1}$ that is not a sphere, projective plane or disc, and $M_{1}$ contains no $2$-sided projective plane as it is a submanifold of $M$ which itself contains no such subsurface. Furthermore, $M_{1}$ is irreducible as any incompressible sphere in $M_{1}$ would represent an incompressible sphere in $M$, and so we may apply Theorem 4.14 to see that $M_{1}\cong\Sigma\times I$. This implies that $M_{1}$ has exactly one other boundary component $\Sigma^{\prime}$ that is isotopic to $\Sigma$, and we therefore have that $\Sigma$ is isotopic to $\Sigma^{\prime}$ in $M$. This gives the desired contradiction, as $\Sigma$ is not isotopic to any boundary component of $M$ by hypothesis.
@@
Step 1 : The action of $G$ on $T$ has exactly one orbit of edges by assumption, so we may apply Theorem 4.6 to see that there exists a non-empty system $\overline{\Sigma}=\{\Sigma_{1},...,\Sigma_{n}\}$ of compact $2$-sided incompressible surfaces embedded in $M$ none of which are boundary parallel such that for all $i$, $\operatorname{Im}(\pi_{1}(\Sigma_{i})\rightarrow\pi_{1}(M))$ is contained in some edge group of $(\Gamma,\mathfrak{G})$ and for all connected components $M_{j}$ of $M\backslash\Sigma$, $\operatorname{Im}(\pi_{1}(M_{j})\rightarrow\pi_{1}(M))$ is contained in some vertex group of $(\Gamma,\mathfrak{G})$. Furthermore, since the boundary components of $M$ are tori, their subgroups are copies of $\mathbb{Z}^{2}$ and so must act elliptically on $T(\Gamma,\mathfrak{G})$ by Lemma 2.19 using the acylindricity assumption. Thus by the second part of the Theorem 4.6 we can assume that all of our $\Sigma_{i}$’s are disjoint from the boundary of $M$, and therefore closed.
Assume first that $M$ has boundary, so any finite sheeted cover of $M$ must also have boundary. Thus, if some cover $M^{\prime}$ of $M$ fibres it must fibre over a surface with at least one boundary component. All virtual surface fibre subgroups of $G$ are therefore the fundamental groups of orientable surfaces with boundary, so must be free groups as such surfaces have free fundamental group. However, each surface in $\Sigma$ may be assumed to be closed by the previous paragraph, and these surfaces are incompressible so cannot be spheres by Theorem 4.8, so the fundamental group of any surface in $\overline{\Sigma}$ will be a finitely generated (from the fact that each surface is compact) non-free group. Thus no surface in $\overline{\Sigma}$ has a fundamental group that includes into $G$ as a virtual surface fibre subgroup, and so they must all be geometrically finite by Theorem 4.12. The case where $M$ has non-empty boundary follows, and we will therefore assume for the remainder of this proof that $M$ is closed.
Step 2 : Assume for contradiction that for all $i\in\{1,...,n\}$ the subgroup $\pi_{1}(\Sigma_{i})$ is a virtual surface fibre subgroup of $G$, and fix some $i$. By construction $\pi_{1}(\Sigma_{i})$ will be contained in some edge stabiliser of the action of $G$ on $T$, which will be some conjugate $G_{e}^{\prime}$ of $G_{e}$ where $G_{e}$ is the unique edge of $\Gamma$. The subgroup $G_{e}^{\prime}$ is an edge stabiliser in a non-elementary and thus non-trivial splitting, so cannot have finite index in $G$, and therefore we may invoke Lemma 4.15 to see that $G_{e}^{\prime}$ is a virtual surface fibre subgroup of $G$, and then $G_{e}$ is conjugate to $G_{e}^{\prime}$ in $G$, so it must also be a virtual surface fibre subgroup.
@@
It follows that our original assumption that for $i\in\{1,..,n\}$, $\pi_{1}(\Sigma_{i})$ is a virtual surface fibre subgroup of $G$ was false, and so by Theorem 4.12 (subgroup tameness) and the fact that each $\pi_{1}(\Sigma_{i})$ is the fundamental group of a closed surface and thus finitely generated, there must exist $i$ such that $\pi_{1}(\Sigma_{i})$ is a geometrically finite subgroup. Thus $M$ must contain an embedded $2$-sided incompressible closed geometrically finite subsurface that is not boundary parallel as required.∎
@@
In this section we use Theorems 1.6 and 4.16 to prove Theorem 1.8, and then briefly show that a similar strong classification result cannot extend to the case where we allow boundary. Theorem 1.8 is restated as follows. See 1.8

### v6

@@ 4.2.1. $I$-bundles
We will require some results about a special type of 3-manifold that is geometrically “nice” in terms of its fundamental group. We will briefly define the concept of an $I$-bundle over a surface, and include a strong result pertaining to $3$-manifolds whose boundary components induce finite index subgroups of the fundamental group. For proof and further exposition see [undefr, Chapter 10].
@@
[undefr, Theorem 10.5] Let $M$ be a compact 3-manifold which contains no 2-sided real projective plane and let $S$ be a compact, connected, incompressible surface embedded in $\partial M$ such that $S$ is not homeomorphic to the disc $B^{2}$, the sphere $S^{2}$ or the projective plane $P^{2}$. If the index $\pi_{1}(M)\colon i_{*}\pi_{1}(S)]$ is finite then either:
@@
    $\pi_{1}(M)\colon\pi_{1}(S)]=1$ and $M=S\times I$ with $S=S\times\{0\}$;
@@
    $\pi_{1}(M)\colon\pi_{1}(S)]=2$ and $M$ is an $I$-bundle over a compact surface $\overline{S}$, with $S$ a two sheeted cover of $\overline{S}$. We call such an $I$-bundle twisted.
@@
All manifolds we consider in this paper will be orientable, and thus cannot contain 2-sided real projective planes due to the fact that, if $P^{2}$ is the real projective plane, $P^{2}\times[-1,1]$ is non-orientable. The condition that $M$ has no two-sided $P^{2}$ has however been included in the statement of Theorem 4.16 for completeness.
@@ 4.3. A Geometric Condition for Acylindrical Arboreality
Let $H$ be a non-trivial virtual surface fibre subgroup of $G$. Then by definition $H$ is finitely generated, and there exists $t\in G$ such that $F=H\rtimes\langle t\rangle$ is a finite index subgroup of $G$ corresponding to a finite sheeted cover of $M$ in which $H$ is a surface fibre subgroup. Let $A$ be a subgroup of $G$ that contains $H$. We separate into two cases.
Case 1, $F\cap A=H$ : In this case, $H$ has finite index in $A$ and is finitely generated, implying that $A$ is also finitely generated. Assume for contradiction that $A$ is not a virtual surface fibre subgroup of $G$. It then follows by Theorem 4.14 that $A$ must be geometrically finite as it is finitely generated, and so by Theorems 4.13 and 4.3 we have that $A$ has finite relative height in $G$ with respect to the boundary subgroups of $G$, which are trivial as $M$ is closed. For all $0\neq j\in\mathbb{Z}$ we have that $t^{j}\notin A$, so the set of conjugates $\{t^{j}At^{-j}\}_{j\in\mathbb{\mathbb{Z}}}$ is essentially distinct, and the intersection of all of these subgroup will contain $H$ as every power of $t$ fixes $H$ by conjugation. Thus, $H$ must be a peripheral or finite subgroup of $G$ by the fact that $A$ has finite relative height, and all of our peripheral subgroups are trivial in this case, so $H$ must be finite. The manifold $M$ is orientable with empty boundary and irreducible by Theorem 4.10, and $G$ admits a non-trivial splitting so it must be infinite, so we may apply Lemma 4.12 to see that $G$ is torsion free, and $H$ in this case must be trivial. However, by Corollary 4.11, $\pi_{1}(M)$ admits a non-trivial splitting over the free subgroup if and only if $M$ is a solid torus, which contradicts the fact that $M$ is closed. It follows that $H$ cannot be trivial, leading to the desired contradiction. Thus, in the case where $A\cap F=H$ we must have that $A$ is a virtual surface fibre subgroup.
Case 2, $A\cap F$ contains $H$ as a proper subgroup : Then there exist $0\neq l\in\mathbb{Z}$ and $h\in H$ such that $t^{l}h\in A\cap F$, but $h\in A\cap F$ by assumption that $A$ contains $H$, and so we must have that $t^{l}\in A\cap F$. It follows that $A\cap F$ contains the subgroup $\langle t^{l},H\rangle$, which has finite index in $H\rtimes\langle t\rangle$, which in turn has finite index in $G$. Thus in this case $A$ contains the finite index subgroup $\langle t^{l},H\rangle$ of $G$, and so $A$ must have finite index in $G$. Thus in both cases either $A$ is a virtual surface fibre subgroup of $G$ or $A$ has finite index in $G$, as required. ∎
@@
Case 1, $\Gamma$ has 2 vertices : In this case we can express $G$ as a non-trivial amalgam $A*_{H}B$ of two subgroups $A$ and $B$ along $H$. In this case, we claim that the subgroup $H$ must have finite index in both $A$ and $B$. Indeed, we first observe that as the amalgam decomposition $A*_{H}B$ is non-trivial by hypothesis on $H$, neither $A$ nor $B$ can have finite index in $G$. Therefore both $A$ and $B$ must be virtual surface fibre subgroups of $G$ by Lemma 4.17, so the normalisers $N_{G}(A)$ and $N_{G}(B)$ of $A$ and $B$ in $G$ must have finite index in $G$. By elementary properties of amalgamated free products and non-triviality of the amalgam $A*_{H}B$, the normaliser $N_{G}(A)$ of $A$ in $G$ must have that $N_{G}(A)\cap B=H$, so it follows from the fact that $N_{G}(A)$ has finite index in $G$ that $H=N_{G}(A)\cap B$ must also be finite index in $B$. Similarly $H$ will have finite index in $A$, and the claim is proved.
The subgroup $H$ is a virtual surface fibre subgroup so it must be isomorphic to the fundamental group of a closed surface, and $M$ is irreducible by Theorem 4.10, so we may apply Theorem 4.7 to realise $H$ geometrically as the fundamental group of some incompressible surface $\Sigma$ in $M$. The manifold $M$ cut along $\Sigma$ then has two components $M_{1}$, with $\pi_{1}(M_{1})=A$, and $M_{2}$, with $\pi_{1}(M_{2})=B$, both 3-manifolds whose fundamental groups are virtually $\pi_{1}(\Sigma)$. In particular, $\Sigma$ must be $2$-sided as it separates $M$, so by the remark after Theorem 4.16 $\Sigma$ is not a real projective plane, and more generally neither $M_{1}$ nor $M_{2}$ can contain a $2$-sided projective plane by the same remark. Furthermore, if $\Sigma$ were a disc we would be able to conclude by Corollary 4.11 that $M$ were a solid torus, contradicting the fact that $M$ is closed, and if $\Sigma$ were a sphere then Theorem 4.10 would imply that $\Sigma$ bounds a ball, contradicting incompressibility. Thus $\Sigma$ is not a disc, sphere, or real projective plane. Therefore, using again the fact that $M$ is irreducible and so each of $M_{1}$ and $M_{2}$ must be irreducible, we may invoke theorem 4.16 on $M_{1}$ (and respectively on $M_{2}$) to see that one of the following occurs.
@@
    $\pi_{1}(M_{1}):\pi_{1}(\Sigma)]=1$, which cannot occur as our original splitting was non-trivial.
@@
    $\pi_{1}(M_{1}):\pi_{1}(\Sigma)]=2$ and $M_{1}$ is a twisted $I$-bundle with boundary $\Sigma$.
@@
Case 2, $\Gamma$ has exactly $1$ vertex : Call this unique vertex $v$ $v$. We can then express $G$ as an HNN extension of $G_{v}$ along $H$, so call the stable letter of this extension $\gamma$. We claim that in this case the subgroup $G_{v}$ must be equal to $H$, and must be a genuine surface fibre subgroup of $G$, i.e. there exists some fibration of $M$ over the circle whose fibre is a surface with fundamental group $G_{v}=H$. Indeed, we must have that either $G_{v}$ is either finite index in $G$ or that $G_{v}$ is a virtual surface fibre subgroup of $G$ by Lemma 4.17. Since $G$ is an HNN extension with base group $G_{v}$, the subgroup $G_{v}$ cannot have finite index, and so we must in fact have that $G_{v}$ is a virtual surface fibre subgroup of $G$. It follows that there exists $t_{v}\in G$ such that $F_{v}=G_{v}\rtimes\langle t_{v}\rangle$ is a finite index subgroup of $G$ corresponding to a cover of $M$ in which $G_{v}$ is a surface fibre subgroup.
Thus there exists some integer $l>0$ such that $\gamma^{l}\in F_{v}$, and so both $\gamma^{l}$ and $\gamma^{-l}$ normalise $G_{v}$. However, by definition of $\gamma$ as the stable letter in the HNN extension of $G_{v}$ along $H$ and by elementary properties of HNN extensions of groups, this can only occur if $H=G_{v}$. It follows that $G=H\rtimes\langle\gamma\rangle$, $H$ is finitely generated by definition of a surface fibre subgroup. Furthermore, $G$ admits a non-trivial splitting and is therefore infinite, so is torsion free by Lemma 4.12, so we have that $H\leq G$ is torsion free and thus $H\not\cong\mathbb{Z}/2\mathbb{Z}$. We may therefore invoke Theorem 4.6 (algebraic fibring) to see that $H=G_{v}$ is a surface fibre subgroup of $G$. The claim is therefore proved, and so in this case $H$ is a surface fibre subgroup so is normal by definition as required.
@@
For the if direction assume that $M$ contains an embedded $2$-sided incompressible geometrically finite closed surface $\Sigma$ that is not isotopic to any boundary component of $M$. Then $\Sigma$ is not a disc as it is closed, $\Sigma$ is not a projective plane by the remark following Theorem 4.16, and $\Sigma$ is not a sphere by Theorem 4.10, so $\pi_{1}(\Sigma)$ (and hence $G$) is infinite by the classification of surfaces (see [undefw, Theorem 4.14], for example). Furthermore, $\pi_{1}(\Sigma)$ is finitely generated as $\Sigma$ is closed. Since $\Sigma$ is geometrically finite we have that $H=\pi_{1}(\Sigma)$ is relatively quasi-convex by Theorem 4.13, and so by Theorem 4.3 there exists a natural number $n_{H}$ such that the intersection of $n_{H}$ essentially distinct conjugates of $H$ in $G$ is either finite or a peripheral subgroup of $G$. In fact, it follows directly from the definition of relatively quasi-convex and by considering the action of $G$ on $\partial\mathbb{H}^{3}$ that $H$ is hyperbolic relative to the subgroups that arise as the intersections of $\Sigma$ with $\partial M$, and that all peripheral elements of $G$ that are included in $H$ are conjugate in $H$ to some boundary component of $\Sigma$. Since $\Sigma$ is closed and therefore does not intersect with $\partial M$, $H$ contains no peripheral elements and so the intersection of any collection of $n_{H}$ essentially distinct conjugates of $H$ in $G$ must be finite. However, $G$ is infinite and $M$ is orientable and irreducible, and so Lemma 4.12 tells us that $G$ and hence $H\leq G$ is torsion free, and thus the intersection of any collection of at least $n_{H}$ essentially distinct conjugates of $H$ in $G$ must be trivial.
@@
Let $(\Gamma,\mathfrak{G})$ be the splitting of $G$ induced by cutting $M$ along $\Sigma$, $T=T((\Gamma,\mathfrak{G})$ its Bass-Serre tree, and let $p$ be a path in $T$ of length at least $n_{H}$. The stabiliser for this path will be the intersection of the stabilisers of its constituent edges, which are essentially distinct conjugates of $H$ in $G$, and so $\operatorname{PStab}_{G}(p)$ is trivial, and the action of $G$ on $T$ is therefore $(n_{H},1)$-acylindrical. The group $G$ is not virtually cyclic as it contains the fundamental group of $\Sigma$, a closed surface that is not a sphere or a projective plane. Thus the action of $G$ on $T$ is not lineal by Theorem 2.3, and to show that it is non-elementary it only remains to check that it is non-trivial.
@@
Assume for contradiction that this splitting is trivial, so $G_{v}=G$ for some vertex $v\in V(\Gamma)$. The graph $\Gamma$ has one edge, which corresponds to $\Sigma$, and so $\Gamma$ either has one vertex or two vertices. The former of these cannot occur, else $G$ would be an HNN extension of $G_{v}$, so $G_{v}\neq G$. We may thus assume that the splitting $(\Gamma,\mathfrak{G})$ is an amalgam decomposition $G\cong G_{v}*_{H}A$ for some subgroup $A$, with $\pi_{1}(\Sigma)=H=A$. Since $\Sigma$ is $2$-sided it must locally separate $M$, and so since the cut along $\Sigma$ induces an amalgam we must have that $\Sigma$ separates $M$ into two components $M_{1}$ and $M_{2}$ with $A=\pi_{1}(M_{1})$ and $G_{v}=\pi_{1}(M_{2})$. By assumption we have that $\pi_{1}(\Sigma)=\pi_{1}(M_{1})$, $\Sigma$ is a connected component of $\partial M_{1}$ that is not a sphere, projective plane or disc, and $M_{1}$ contains no $2$-sided projective plane as it is a submanifold of $M$ which itself contains no such subsurface. Furthermore, $M_{1}$ is irreducible as any incompressible sphere in $M_{1}$ would represent an incompressible sphere in $M$, and so we may apply Theorem 4.16 to see that $M_{1}\cong\Sigma\times I$. This implies that $M_{1}$ has exactly one other boundary component $\Sigma^{\prime}$ that is isotopic to $\Sigma$, and we therefore have that $\Sigma$ is isotopic to $\Sigma^{\prime}$ in $M$. This gives the desired contradiction, as $\Sigma$ is not isotopic to any boundary component of $M$ by hypothesis.
@@
Step 1 : The action of $G$ on $T$ has exactly one orbit of edges by assumption, so we may apply Theorem 4.8 to see that there exists a non-empty system
$\overline{\Sigma}=\{\Sigma_{1},...,\Sigma_{n}\}$
of compact $2$-sided incompressible surfaces embedded in $M$ none of which are boundary parallel such that $\operatorname{Im}(\pi_{1}(\Sigma_{i})\rightarrow\pi_{1}(M))$ for all $i$ is contained in some edge group of $(\Gamma,\mathfrak{G})$ and $\operatorname{Im}(\pi_{1}(M_{j})\rightarrow\pi_{1}(M))$ for all connected components $M_{j}$ of $M\backslash\Sigma$ is contained in some vertex group of $(\Gamma,\mathfrak{G})$. Furthermore, since the boundary components of $M$ are tori, their subgroups are copies of $\mathbb{Z}^{2}$ and so must act elliptically on $T(\Gamma,\mathfrak{G})$ by Lemma 2.19 using the acylindricity assumption. Thus by the second part of Theorem 4.8 we can assume that all of our $\Sigma_{i}$’s are disjoint from the boundary of $M$, and therefore closed.
Assume first that $M$ has boundary, so any finite sheeted cover of $M$ must also have boundary. Thus, if some cover $M^{\prime}$ of $M$ fibres it must fibre over a surface with at least one boundary component. All virtual surface fibre subgroups of $G$ are therefore the fundamental groups of orientable surfaces with boundary, so must be free groups as such surfaces admit deformation retractions onto graphs. However, each surface in $\overline{\Sigma}$ may be assumed to be closed by the previous paragraph, and these surfaces are incompressible so cannot be spheres by Theorem 4.10. The fundamental group of any surface in $\overline{\Sigma}$ will therefore be a finitely generated (from the fact that each surface is compact) group that is not a free group owing to the fact that such groups have cohomological dimension $2$ [undeff, Chapter VIII.2, Example 3] but free groups have cohomological dimension $1$ [undeff, Chapter VIII.2, Example 2] (see [undeff, Chapter VIII] for a definition of cohomological dimension). Thus no surface in $\overline{\Sigma}$ has a fundamental group that includes into $G$ as a virtual surface fibre subgroup, and so they must all be geometrically finite by Theorem 4.14. The case where $M$ has non-empty boundary follows, and we will therefore assume for the remainder of this proof that $M$ is closed.
Step 2 : Assume for contradiction that for all $i\in\{1,...,n\}$ the subgroup $\pi_{1}(\Sigma_{i})$ is a virtual surface fibre subgroup of $G$, and fix some $i$. By construction $\pi_{1}(\Sigma_{i})$ will be contained in some edge stabiliser of the action of $G$ on $T$, which will be some conjugate $G_{e}^{\prime}$ of $G_{e}$ where $e$ is the unique edge of $\Gamma$. The subgroup $G_{e}^{\prime}$ is an edge stabiliser in a non-elementary and thus non-trivial splitting, so cannot have finite index in $G$, and therefore we may invoke Lemma 4.17 to see that $G_{e}^{\prime}$ is a virtual surface fibre subgroup of $G$, and then $G_{e}$ is conjugate to $G_{e}^{\prime}$ in $G$, so it must also be a virtual surface fibre subgroup.
@@
It follows that our original assumption that for $i\in\{1,..,n\}$, $\pi_{1}(\Sigma_{i})$ is a virtual surface fibre subgroup of $G$ was false, and so by Theorem 4.14 (subgroup tameness) and the fact that each $\pi_{1}(\Sigma_{i})$ is the fundamental group of a closed surface and thus finitely generated, there must exist $i$ such that $\pi_{1}(\Sigma_{i})$ is a geometrically finite subgroup. The surface $\Sigma_{i}$ was constructed using Theorem 4.8, and so is not boundary parallel, and $M$ must contain an embedded $2$-sided incompressible closed geometrically finite subsurface that is not boundary parallel as required.∎
@@
In this section we use Theorems 1.6 and 4.18 to prove Theorem 1.8, and then briefly show that a similar strong classification result cannot extend to the case where we allow boundary. Theorem 1.8 is restated as follows. See 1.8